# ZigzagNetVis: Suggesting temporal resolutions for graph visualization using zigzag persistence

### v2

@@ 2.3 Zigzag persistent homology
The bottleneck distance lies at the core of our method and will be used as a means to select resolutions in Sec. 4. Namely, we will compare the temporal graphs coming from two different resolutions via the bottleneck distance between the persistence barcodes coming from their zigzag filtrations. This distance computes the global topological agreement between these temporal graphs, allowing us to determine whether they are similar or not, just as in the context of abrupt change detection. In addition, the bottleneck distance offers two advantages. First, it allows for a theoretical treatment: we will study in Sec. 4.2 what values of distance are to be expected, and when they indicate a relevant change. Secondly, and as a consequence of its definition, the bottleneck distance is always caused by a pair of bars or a bar alone. From a practical point of view, one can identify which nodes of the graph are responsible for the topological difference between two barcodes. Based on this observation, we will describe an explainability pipeline in Sec. 9.1.
@@ 4.1 Description of the method
In a nutshell, our methodology employs the bottleneck distance between consecutive resolutions as a feature to perform abrupt change detection. While change detection based on features is common in temporal graph analysis (see, e.g., [3]), incorporating PH offers several advantages. First, thanks to the high interpretability of PH, we can give a heuristic analysis in Sec. 4.2, already yielding important insights. Moreover, as studied further in Sec. 9, the bottleneck distance appears to be a stable and relevant quantity, gathering information from various other features of the literature.
@@ 4.2 Timestamps shifts and structural changes
Peak detection. In real-life examples, the normalized suggestion curve may contain many positive values. However, returning all the corresponding resolutions to the user would not be relevant. Instead, we choose to return only the most prominent peaks of the curve. In practice, prominence is computed using the package signal of scipy. We return only $m=5$ maxima, five being an arbitrary value that we found suitable. We will give concrete outputs of our algorithm on eight temporal graphs in Sec. 9.1.
Other distances. In TDA, one chooses a distance according to the context: while the bottleneck distance calculates the maximal discrepancy between two barcodes, the Wasserstein distance incorporates all perturbations. The latter option is of interest, for instance, when low-persisting features matter [14]. In contrast, our work aims to detect structural changes, which are evidenced by the perturbation of a single bar of the barcode. Therefore, the bottleneck distance appears as a natural choice (see, for instance, Fig. 2). This observation is supported by Sec. C.2.4 of our supp. material, where it is shown that the Wasserstein distance leads to less interpretable results. In the same vein, one could use, instead of the bottleneck distance, any feature that quantifies the proximity between two temporal graphs. To this end, many quantities exist, such as those presented in Sec. 9.2 (e.g., mean degree, density, or burstiness). However, they all appear to either lack stability or provide limited information. Our experimental study shows that the bottleneck distance acts as a relevant trade-off between stability and information, “incorporating” several popular features.
@@ 5.2 Design decisions
Alternative designs. We have considered alternative design choices before proposing this layout. We decided to use a timeline visualization instead of animations to better meet analyses that rely on multiple and often distant timestamps (tasks T2, T3), complying with [5]. We then studied the suitability of existing timeline visualizations for our context. An option would be a visualization based on tracking graphs [34, 66, 37]. In particular, LargeNetVis’s Global View [34] is a grid-based layout where rows and columns represent respectively network communities and timeslices. In this view (see Fig. 11(b)), communities are encoded as circles with varying sizes, and their temporal evolution is depicted through links connecting communities from consecutive timeslices. Although we could adapt it to encode connected components, it would still not provide immediate information about components’ node members. The identification and tracking of the members would also not be immediate with other visualizations, for instance, MSV [61], PAOH [60], and TAM [34].
@@ 7 Usage Scenario
We focus on two types of analysis. First, we demonstrate the suitability of a suggested resolution for analyzing the Primary School and the usefulness of our colored barcode and system to assist in this analysis. Later, we show that the patterns found using ZigzagNetVis are comparable to those identified using LargeNetVis [34], a state-of-the-art approach.
7.1 Exploring the Primary School network
@@
7.2 Comparison with LargeNetVis
To validate the colored barcode, we performed a direct comparison with LargeNetVis [34], an established approach to visualize large temporal networks. To be coherent with the partition timeslicing used by LargeNetVis, we decided to use partition timeslicing in ZigzagNetVis as well (see Sec. B.2 in our supp. material for a visual comparison between partition and sliding timeslicing). We also forced the number of timeslices in LargeNetVis to be equal to the number of partitions in ZigzagNetVis for a fair comparison.
Using the Primary School network, Fig. 11 shows a comparison between our colored barcode and LargeNetVis’ Global View, also showing node-link diagrams that support the comparison. The first highlighted pattern refers to a single connected component on ZigzagNetVis containing students and teachers from three classes (4A, 5A, 5B) (Fig. 11(a,I)) and the two equivalent communities from LargeNetVis (Fig. 11(b,I)). When analyzing the corresponding node-link diagrams (Fig. 11(c)), we see that these two communities form a single connected component thanks to a single edge (dotted in red) linking two teachers (task T1). Also, we see that students from class 4A interact with each other but not with the other two classes (5A and 5B); on the other hand, students from 5A interact with students from 5B and vice-versa (T1). This finding is supported by the fact that students in the same class interact more often with themselves than with students in other classes and that the same goes for same-grade students. Note also that LargeNetVis only allows us to identify the school classes that are found in a community through the node-link diagram. However, this information is immediate with ZigzagNetVis’ colored barcode.
Figure 11: Comparison between (a) ZigzagNetVis with bottom-based ordering and (b) LargeNetVis highlighting three distinct patterns (I-III) in the Primary School network.
Regarding the second pattern (Fig. 11(a-b,II)), both layouts were able to faithfully represent the continuous level of interactions involving students in class 2B and their teacher (task T2). Once again, note that the colored barcode already shows the school class involved in the interactions. Regarding the third pattern, the colored barcode highlights a component that contains students from five different classes interacting with each other during lunch break (Fig. 11(a,III)). When analyzing the node-link diagram (Fig. 11(d)), we see several interactions between these students during that period, i.e., the component is strongly connected (task T1). In LargeNetVis, due to the nature of the community detection algorithm, this strongly connected component was divided into seven small communities, which impaired the finding of this strongly connected group (see Fig. 11(b,III) and Fig. 11(e)).
Each layout has advantages and disadvantages depending on the user task. We aimed to demonstrate that our approach compares to well-validated visualizations, producing equally relevant results.
@@ 8.3 Results
Hypotheses on data analysis. All participants answered at least one of the points that we expected for each open question (SQ1-SQ5, SQ7, SQ8). Also, during the experiment, we encouraged participants to raise hypotheses that could justify specific patterns considering a school environment. For instance, in question SQ1 (Primary School), we asked them to evaluate the relationship between students and teachers from classes 4A, 5A, and 5B (which form a connected component at some point, as shown in Fig. 11(I)). All participants mentioned that class 4A was far from the others. Furthermore, 62% of the participants identified that the two subgroups (4A, 5A-5B) were linked by an edge that involved a teacher; 37% noticed that this edge actually involves two teachers. The hypotheses put forward to explain the strong interaction between classes 5A and 5B mentioned that, since both classes belong to the fifth grade (30% remembered this information), it could be due to interdisciplinary events such as laboratory activity (22%) or group studies (14.81%).
@@ 9 Analysis of the suggested resolutions
This section is devoted to the analysis of the resolutions suggested by ZigzagNetVis’ methodology. We aim to demonstrating empirically that these suggestions are relevant. As mentioned in Sec. 2, few studies have tackled the problem of choosing a resolution. In particular, no reference data sets are available. As a means of comparison, we will use, for each of the temporal graphs considered in this paper, the resolutions used in the literature — that we stress are mainly chosen “by hand”. However, a direct comparison cannot be made. Indeed, as pointed out in Sec. 4.1, one cannot define “optimal” resolutions, but rather meaningful ranges of resolutions. Consequently, to assess the quality of ZigzagNetVis’ suggested resolutions, one must analyze them qualitatively by understanding the behaviors of the corresponding dynamic graphs and comparing them with the literature. This study will be conducted in Sec. 9.1, in particular using the visual tool provided by the bottleneck distance. In Sec. 9.2, we extend this study by comparing our suggestion curves with commonly used feature of temporal graphs, revealing when they coincide and when they complement each other.
@@ 9.1 Visual justification with colored barcodes
This classification allows, at least heuristically, to divide the range of resolutions into three intervals: a chaotic start, followed by a range where the resolution curve only exhibits relevant peaks, and at last a few values caused by the merging of the days. This observation can be used when the user, through manual inspection, seeks relevant resolutions to study. We stress that the last case is not observed in the figures the Primary and High School networks, since we selected resolutions smaller than the length of a night.
9.2 Comparison with other features
We now compare our novel method with existing techniques. In the literature, features of temporal graphs are of two sorts: either they are features of (non-temporal) graphs, adapted to the temporal case by taking their mean of their list over all the snapshots, or they directly depend on the temporal structure [54, 44].
Geometric and topological features of snapshots. Let $G$ be a temporal graph to which we apply a sliding-window timeslicing of resolution $r$. Given a snapshot $G_{t}$, we consider:
  • its number of nodes and edges, denoted $N(t)$ and $E(t)$,
  • its density $D(t)=2E(t)/(V(t)(V(t)-1))$,
  • its number of connected components $CC(t)$,
  • the mean degree $MD(t)$ of its nodes,
  • its transitivity $T(t)$, defined the ratio between the number of triangles and triads (i.e., pairs of edges sharing a vertex).
Taking the mean value of such a feature over all times $t$ yields a feature of the temporal graph $G$. We denote them respectively $N(r)$, $E(r)$, $D(r)$, $CC(r)$, $MD(r)$ and $T(r)$, making explicit the dependence on the resolution.
Note that, when increasing the resolution $r$, the features $N(r)$, $E(r)$ and $D(r)$ increase. That is, they are non-decreasing functions. In general, we except that abrupt changes in these values reflect the fact that the temporal graph exhibits a new behavior. To visualize such changes, one can plot these curves or, more efficiently, their derivative. These curves are represented in Fig. 17 for the Primary and High School networks, and in Fig. 8 of our supp. material for the other graphs. Since we are only interested in the peaks or qualitative behaviors of these curves, and not their absolute values, we normalize them so that their maxima equals one. In order to ease the reading, the $x$-axis is divided in two windows, and the curves are normalized both times.
Figure 17: Suggestion curve and derivative of the mean snapshot features for the networks Primary School (top) and High School (bottom). The $x$ axis represents the resolution values. The curves are normalized, and a few interesting values are highlighted with a dashed line.
A manual inspection of these curves allows us to compare them with our suggestion curve. For instance, in the Primary School Network, one sees that the first peak, at $r=8$, corresponds to the global maximum of the derivative of the number of connected components. Besides, the peak at resolution 154 seems to appear simultaneously as the transitivity curve shows a constant derivative.
Similar observations can be made on the High School network. The peaks of our suggestion curve found at resolutions 46, 204, and 216 correspond, respectively, to global maxima of the derivative of the number of connected components, transitivity, and density. Besides, the peaks at resolutions 12 and 56 correspond, respectively, to a global minimum of the derivative of the number of connected components and a significant local maximum of the transitivity.
These observations suggest that our curve captures information coming from various features of graphs. However, some peaks remain unexplained, and hence we will study other features in the paragraphs below. It must also be noted that certain features’ peaks do not correspond to a peak of the suggestion curve. This may be caused by the fact that the suggestion curve, based on the homology group $H_{0}$, is blind to certain purely geometric properties of graphs, and works only in terms of connected components.
Distribution of features of snapshots. Instead of taking the average value of a feature over all the snapshots, we can compare their distribution over time. To do so, we consider the entire curves
$f_{N,r}\colon t\mapsto N(t)$,     $f_{E,r}\colon t\mapsto E(t)$,     etc.
Given two consecutive resolutions $r$ and $r+2$, we compare these curves via their $\ell^{2}$-norm
$\|f_{N,r}-f_{N,r+2}\|_{2}$,     $\|f_{E,r}-f_{E,r+2}\|_{2}$,     etc.
These are functions of $r$, on which we expect to observe abrupt changes in the graph’s behavior. These curves are represented in Fig. 18 for the Primary and High School networks and in Fig. 9 of our supp. material for the other graphs.
Figure 18: Consecutive distances between the distribution of the snapshot features for the networks Primary School (top) and High School (bottom). The $x$ axis represents the resolution values. The curves are normalized, and a few interesting values are highlighted with a dashed line.
As before, one draws correspondences between the suggestion curve and these features. For example, on the Primary School network, the first suggested resolution, $r=8$, happens precisely during an abrupt change in the derivative of all the curves. Besides, the two peaks at resolutions 120 and 154 delimit the only interval where the curve of transitivity increases and then decreases. This last correspondence has already been observed in the last paragraph while considering the mean transitivity of the temporal graph.
In a similar fashion, on the High School network, one observes that the selected resolution 46 corresponds to an abrupt change in the curve built from transitivity. In a few words, comparing the distribution of the snapshot features offers compatible but also complementary information to the average values alone.
Global features. Lastly, we consider features of dynamic graphs that do not come from features of snapshots. Given a temporal graph $G$, timesliced at a resolution $r$, we consider:
  • its burstiness $B(r)$ and average lifecycle $LC(r)$, defined in [53].
  • its stability $S(r)$ and fidelity $F(r)$, defined in [15].
Moreover, we will also consider the total persistence $TP(r)$ of its corresponding zigzag persistence module, defined as the quadratic mean of the length of its bars. Finally, we will employ the curve $MDS(r)$ defined in [25]. It consists of the multidimensional scaling (MDS) in dimension 1, whose input is the set of bottleneck distances between the persistence barcodes of the temporal graph for all the resolutions considered. These curves are represented in Fig. 19 for the Primary and High School networks and in Fig. 10 of our supp. material for the other graphs.
Figure 19: Global features as functions of the resolution parameter, for the networks Primary School (top) and High School (bottom). The $x$ axis represents the resolution values. The curves are normalized, and a few interesting values are highlighted with a dashed line.
On all the graphs, one observed a high correlation between our suggestion curve and the curves of MDS and total persistence. This is expected since all these features are related to the persistence barcodes of the zigzag modules. We also observe that the peak at $r=18$ of our suggestion curve for the Primary School corresponds to a global minimum of the lifecycle. The same occurs with $r=12$ for the High School network.
Regarding stability and fidelity, we observe that the curves $r\mapsto S(r)$ and $r\mapsto F(r)$ are convex and no abrupt change can be observed. Instead, we consider a relevant feature that can be defined from them: the point of intersection between the normalized stability and the inverse of the normalized fidelity reveals the resolution value that balances the best between these antagonists’ values. The resolutions suggested by this strategy are $r=24$ and $r=12$ for the Primary and High School, respectively. Note that ZigzagNetVis also suggests $r=12$ for the High School.
In conclusion of this section, the peaks of the suggestion curve can, most of the time, be mapped to peaks or bumps of other features in the literature. We stress that our analysis does not reveal the exact nature of this connection; we simply showed how the suggestion curve can be understood as related to other features.
@@ 10 Discussion and Limitations
Resolution comparison. Some participants would also like to simultaneously compare suggested resolutions with each other and with non-suggested ones. Although one could open the system many times or perform a side-by-side comparison using multiple instances of the system, we believe that incorporating such a capability into our system would enhance the identification of patterns coming from different resolutions, help users to understand and follow changes that regions of interest suffer when varying resolutions (recall Sec. 9.1), and also increase the user’s confidence in the suggestions or reveal room for improvement in the suggestion procedure, e.g., by incorporating user feedback.

### v3

@@ 2.3 Zigzag persistent homology
The bottleneck distance lies at the core of our method and will be used as a means to select resolutions in Sec. 4. Namely, we will compare the temporal graphs coming from two different resolutions via the bottleneck distance between the persistence barcodes coming from their zigzag filtrations. This distance computes the global topological agreement between these temporal graphs, allowing us to determine whether they are similar or not, just as in the context of abrupt change detection. In addition, the bottleneck distance offers two advantages. First, it allows for a theoretical treatment: we will study in Sec. 4.2 what values of distance are to be expected, and when they indicate a relevant change. Secondly, and as a consequence of its definition, the bottleneck distance is always caused by a pair of bars or a bar alone. From a practical point of view, one can identify which nodes of the graph are responsible for the topological difference between two barcodes. Based on this observation, we will describe an explainability pipeline in Sec. 9.
@@ 4.1 Description of the method
In a nutshell, our methodology employs the bottleneck distance between consecutive resolutions as a feature to perform abrupt change detection. While change detection based on features is common in temporal graph analysis (see, e.g., [3]), incorporating PH offers several advantages. First, thanks to the high interpretability of PH, we can give a heuristic analysis in Sec. 4.2, already yielding important insights. Moreover, as studied further in the supp. material, the bottleneck distance appears to be a stable and relevant quantity, gathering information from various other features of the literature.
@@ 4.2 Timestamps shifts and structural changes
Peak detection. In real-life examples, the normalized suggestion curve may contain many positive values. However, returning all the corresponding resolutions to the user would not be relevant. Instead, we choose to return only the most prominent peaks of the curve. In practice, prominence is computed using the package signal of scipy. We return only $m=5$ maxima, five being an arbitrary value that we found suitable. We will give concrete outputs of our algorithm on eight temporal graphs in Section 9.
Other distances. In TDA, one chooses a distance according to the context: while the bottleneck distance calculates the maximal discrepancy between two barcodes, the Wasserstein distance incorporates all perturbations. The latter option is of interest, for instance, when low-persisting features matter [14]. In contrast, our work aims to detect structural changes, which are evidenced by the perturbation of a single bar of the barcode. Therefore, the bottleneck distance appears as a natural choice (see, for instance, Fig. 2). This observation is supported by Sec. C.2.4 of our supp. material, where it is shown that the Wasserstein distance leads to less interpretable results. In the same vein, one could use, instead of the bottleneck distance, any feature that quantifies the proximity between two temporal graphs. To this end, many quantities exist, such as those presented in Section C.2.3 of our supp. material (e.g., mean degree, density, or burstiness). However, they all appear to either lack stability or provide limited information. Our experimental study shows that the bottleneck distance acts as a relevant trade-off between stability and information, “incorporating” several popular features.
@@ 5.2 Design decisions
Alternative designs. We have considered alternative design choices before proposing this layout. We decided to use a timeline visualization instead of animations to better meet analyses that rely on multiple and often distant timestamps (tasks T2, T3), complying with [5]. We then studied the suitability of existing timeline visualizations for our context. An option would be a visualization based on tracking graphs [34, 65, 37]. In particular, LargeNetVis’s Global View [34] is a grid-based layout where rows and columns represent respectively network communities and timeslices. In this view, communities are encoded as circles with varying sizes, and their temporal evolution is depicted through links connecting communities from consecutive timeslices. Although we could adapt it to encode connected components, it would still not provide immediate information about components’ node members. The identification and tracking of the members would also not be immediate with other visualizations, for instance, MSV [60], PAOH [59], and TAM [34].
@@ 7 Usage Scenario
We demonstrate the suitability of a suggested resolution for analyzing the Primary School and the usefulness of our colored barcode and system to assist in this analysis. We also show in Section C.4 of our supp. material that the patterns found using ZigzagNetVis are comparable to those identified using LargeNetVis [34], a state-of-the-art approach.
@@ 8.3 Results
Hypotheses on data analysis. All participants answered at least one of the points that we expected for each open question (SQ1-SQ5, SQ7, SQ8). Also, during the experiment, we encouraged participants to raise hypotheses that could justify specific patterns considering a school environment. For instance, in question SQ1 (Primary School), we asked them to evaluate the relationship between students and teachers from classes 4A, 5A, and 5B (which form a connected component at some point). All participants mentioned that class 4A was far from the others. Furthermore, 62% of the participants identified that the two subgroups (4A, 5A-5B) were linked by an edge that involved a teacher; 37% noticed that this edge actually involves two teachers. The hypotheses put forward to explain the strong interaction between classes 5A and 5B mentioned that, since both classes belong to the fifth grade (30% remembered this information), it could be due to interdisciplinary events such as laboratory activity (22%) or group studies (14.81%).
@@ 9 Analysis of the suggested resolutions
This section is devoted to the analysis of the resolutions suggested by ZigzagNetVis’ methodology. We aim to demonstrating empirically that these suggestions are relevant. As mentioned in Sec. 2, few studies have tackled the problem of choosing a resolution. In particular, no reference data sets are available. As a means of comparison, we will use, for each of the temporal graphs considered in this paper, the resolutions used in the literature — that we stress are mainly chosen “by hand”. However, a direct comparison cannot be made. Indeed, as pointed out in Sec. 4.1, one cannot define “optimal” resolutions, but rather meaningful ranges of resolutions. Consequently, to assess the quality of ZigzagNetVis’ suggested resolutions, one must analyze them qualitatively by understanding the behaviors of the corresponding dynamic graphs and comparing them with the literature. This study will be conducted here, in particular using the visual tool provided by the bottleneck distance. In our supp. material (Section C.2.3), we extend this study by comparing our suggestion curves with commonly used feature of temporal graphs, revealing when they coincide and when they complement each other.
@@
This classification allows, at least heuristically, to divide the range of resolutions into three intervals: a chaotic start, followed by a range where the resolution curve only exhibits relevant peaks, and at last a few values caused by the merging of the days. This observation can be used when the user, through manual inspection, seeks relevant resolutions to study. We stress that the last case is not observed in the figures for the Primary and High School networks, since we selected resolutions smaller than the length of a night.
@@ 10 Discussion and Limitations
Resolution comparison. Some participants would also like to simultaneously compare suggested resolutions with each other and with non-suggested ones. Although one could open the system many times or perform a side-by-side comparison using multiple instances of the system, we believe that incorporating such a capability into our system would enhance the identification of patterns coming from different resolutions, help users to understand and follow changes that regions of interest suffer when varying resolutions (recall Sec. 9), and also increase the user’s confidence in the suggestions or reveal room for improvement in the suggestion procedure, e.g., by incorporating user feedback.